OpenStudy (anonymous):

what is the slope of a line that runs parallel to y = 2x + 5

Vocaloid (vocaloid):

well, do you know how to find the slope of y = 2x + 5?

OpenStudy (anonymous):

no can you help me? sorry i know stupid question

Vocaloid (vocaloid):

nah, it's fine, I'm happy to help the slope-intercept form of a line is y = mx + b, where m is the slope so let's put our two equations side by side for comparison y = mx + b y = 2x + 5 can you tell me what number matches with m?

OpenStudy (anonymous):

2

Vocaloid (vocaloid):

yup! so our slope is 2 now, a parallel line will have the same slope, so the new line will also have a slope of 2 so 2 is our final answer

OpenStudy (anonymous):

okay thank you soo muchh!!! hahha that was easy! may i ask for help for a two more??

Vocaloid (vocaloid):

sure

OpenStudy (anonymous):

A rectangle with a length of x - 4 and a width of 8 has a perimeter of 34. What is the value of x?

Vocaloid (vocaloid):

well, the perimeter of a rectangle is 2(width+length) so we can take that information and write an equation 34 = 2(x-4+8) do you know how to find the value of x from here?

OpenStudy (anonymous):

will i be doing cross multiply? or is that something else?

OpenStudy (anonymous):

@Vocaloid

Vocaloid (vocaloid):

not quite, cross multiplying is when we have fractions I'll walk you through it so, we start off with 34 = 2(x-4+8) we can divide both sides by 2 to get 17 = (x-4+8) with me so far?

OpenStudy (anonymous):

yes

Vocaloid (vocaloid):

alright, now we can clean up the inside of the parentheses 17 = (x-4+8) 17 = x + 4 can you tell me what x is now?

OpenStudy (anonymous):

will i devide the 4? same?

Vocaloid (vocaloid):

almost but not quite the opposite of addition is subtraction, correct? since we have 4 added to one side, we can subtract 4 from both sides to get rid of it 17 - 4 = x + 4 - 4 13 = x

OpenStudy (anonymous):

yes yes ohh okay i got it now!! ahh thank thank you soo much! i feel so stupid but thank you for patience. one more please help one more?

Vocaloid (vocaloid):

sure, and don't apologize, you're here to learn :)

OpenStudy (anonymous):

What is the solution to 4x -14 + 6 = 12? thank you so much

Vocaloid (vocaloid):

well, let's start by looking at the left side we can put -14 and 6 together since these are both constants (no letters attached to them( -14 + 6 = ?

OpenStudy (anonymous):

-8

Vocaloid (vocaloid):

good, so we put them together to get -8 now we have 4x - 8 = 12 now, since we have minus 8 on one side, we can +8 to both sides to get rid of the 8 so 12 + 8 = ?

OpenStudy (anonymous):

20

Vocaloid (vocaloid):

yup, so now we've got 4x = 20 now divide both sides by 4, and what do you get?

OpenStudy (anonymous):

x=5

OpenStudy (anonymous):

right? im so happy im learning jeje

Vocaloid (vocaloid):

yup, excellent work

OpenStudy (anonymous):

than thank you so much @vocaloid!!